What happens when Prussian Blue (KFe[Fe(CN)6]) is added to the three solutions (3M NaOH, 3M HCl, 10% H2SO4) at room temperature? Please include the chemical equations.
1
Expert's answer
2017-03-26T05:00:07-0400
What happens when Prussian Blue (KFe[Fe(CN)6]) is added to the three solutions (3M NaOH, 3M HCl, 10% H2SO4) at room temperature? Please include the chemical equations.

K3[Fe(CN)6] + 3 NaOH = Fe(OH)3 + 3 NaCN + 3 KCN
FeSO4 + K3[Fe(CN)6] + 2 HCl = KFe[Fe(CN)6] + H2SO4 + 2 KCl
K3[Fe(CN)6] + 6H2SO4 + 6H2O = 2K2SO4 + FeSO4 + 3(NH4)2SO4 + 6CO